1) A disk with a c value of 1/2, a mass of 4 kg, and radius of 0.54 meters, rolls without slipping down an incline with has a length of 7 meters and angle of 30 degrees. At the top of the incline the disk is spinning at 24 rad/s. How fast is the it moving (the center of mass) at the top of the incline in m/s?
Answer #1

w = 24 radis r=o.sum - gsino -- using W2=42+295 V = V2 + 2 gsine x5 Vi= swi= 0.54X24 = 12-36 m/s S=7m 9=300 sinso = 1 V2 = (1